OpenStudy (anonymous):

So... ya I have no idea. HELP (2x+6)-4=20 x=9 and x=11 x=-9 and x=15 x=9 and x=-15

OpenStudy (anonymous):

What i got was X=9.......

OpenStudy (anonymous):

You need to isolate the variable (2x) on one side of the equation to solve.

OpenStudy (anonymous):

Yes you are correct about x=9 but what do you think the other number is?

OpenStudy (anonymous):

I do not know how to get the other number that is what i need help with.

OpenStudy (anonymous):

You sure that it is not suppose to be |2x+6| -4 =20

OpenStudy (anonymous):

it is |2x+6| -4 =20 what did i put that is different

OpenStudy (anonymous):

Alright so an easy way of figuring it out is by plugging in the numbers so for example for this equation you would take (2(11)+6)-4=20. Then you would get 24=20 which would be incorrect so then you would take the next number and plug it in.

OpenStudy (anonymous):

Parenthesis..to solve for absolute value you would solve the equation as shown but solve for 20 and -20. ie 2x+6-4 = 20 and 2x+6-4 = -20

OpenStudy (anonymous):

but when i do -20 i get -11 that is not a answer

OpenStudy (anonymous):

@nelsonjedi